# Learning Macroeconomic Policies based on Microfoundations: A Stackelberg Mean Field Game Approach

### v1

@@ 5 Experiment
Figure 2: The training curves for 6 algorithms on 4 macroeconomic indicators, comparing settings without behavior cloning as pre-train (N=100 without BC) and with behavior cloning (N=100-BC & N=1000-BC).
Table 1: Test results for 6 algorithms on 5 indicators under 3 settings (N=100 without BC / N=100-BC / N=1000-BC).
@@
Algorithms Per Capita GDP Social Welfare Income Gini Wealth Gini Years
Rule-based 1.41e+05 / 3.66e+11 / 1.41e+05 69.27 / 79.23 / 334.79 0.89 / 0.52 / 0.90 0.92 / 0.53 / 0.93 1.00 / 217.45 / 1.00
DDPG 1.41e+05 / 2.03e+12 / 4.92e+11 70.91 / 94.50 / 527.09 0.88 / 0.46 / 0.75 0.92 / 0.48 / 0.79 1.00 / 299.85 / 100.68
MADDPG 4.93e+04 / 6.38e+12 / 6.82e+12 55.74 / 93.89 / 954.88 0.91 / 0.57 / 0.56 0.92 / 0.58 / 0.62 1.00 / 268.53 / 278.50
IDDPG 1.21e+05 / 7.41e+12 / 2.79e+12 83.09 / 98.16 / 512.19 0.88 / 0.53 / 0.77 0.92 / 0.55 / 0.81 1.00 / 300.00 / 100.68
MF-MARL 8.66e+07 / 5.44e+12 / 1.13e+05 82.02 / 98.21/ 440.00 0.82 / 0.50 / 0.90 0.83 / 0.52 / 0.93 75.75 / 300.00 / 1.00
SMFRL 9.59e+12 / 1.01e+13 / 9.68e+12 96.87 / 96.90 / 975.15 0.52 / 0.51 / 0.52 0.51 / 0.53 / 0.51 300.00 / 299.89 / 300.00
In this section, we conduct two important experiments in the TaxAI environment: (1) We compare the performance of the SMFRL algorithm with 5 baselines in solving SMFGs; (2) We compare the effects of the SMFG method with 5 economic policies, and discuss the efficiency-equity tradeoff and SMFG assumptions.
5.1 Environment
TaxAI (Mi et al., 2023) is a reinforcement learning simulator used for studying optimal tax policy. The environmental settings include the government promoting economic growth (i.e., GDP) through adjusting tax policy and government spending, and households optimizing labor supply and savings ratio for long-term personal utility. It is one of the few simulators for researching optimal macroeconomic policies, hence we conduct experiments on it.
5.2 Algorithmic Analysis for SMFGs
Baselines
We compare popular reinforcement learning algorithms including DDPG (Lillicrap et al., 2015) (only for leader), IDDPG (de Witt et al., 2020), MADDPG (Lowe et al., 2017), Mean Field MARL (Yang et al., 2018), and rule-based policy with our algorithm SMFRL. For more details see Appendix B.1 and Table 2.
Results
We compared SMFRL with 5 baselines in three different experimental setups: without behavior cloning as pre-training and with 100 followers (marked as N=100 without BC); with BC-based pre-training for follower agents at N=100 and N=1000 (N=100-BC; N=1000-BC). Figure 2 illustrates the training curves of 4 key macroeconomic indicators under these three settings. Each row corresponds to one setting, and each column to a macroeconomic indicator, including per capita GDP, social welfare, income Gini, and wealth Gini. A rise in per capita GDP indicates economic growth, an increase in social welfare implies happier households and a lower Gini index indicates a fairer society. Each subplot’s Y-axis represents the indicators’ values, and the X-axis represents the training steps. Table 1 displays the test results of the 6 algorithms across 5 indicators, with each column corresponding to an experimental setting.
5.3 Comparison of Macroeconomic Policies
Baselines
We compare free market policy, Saez tax (Saez, 2001), U.S. Federal tax  [^2] https://data.oecd.org/, AI Economist (Zheng et al., ), and AI Economist-BC. Saez tax will provide suggestions for real-world tax reform. For more details see Appendix B.1.
Policies Results
Figure 5: Comparison of micro and macroeconomic indicators at varying proportions of SMFG households, demonstrating the positive impact on the individual and overall economy of SMFG households.
Efficiency-Equity Tradeoff
SMFG Assumption Analysis
The above experiments have already demonstrated the superiority of the SMFG method. However, the modeling and solving of SMFG rely on an important assumption: the optimal decisions of governments and households depend on the overall state information. In reality, it is natural for the government to focus on market information in decision-making, but it cannot be guaranteed that all households will consider market information in their decisions. To discuss the applicability of the SMFG method, we introduce two types of households within the same economy: those based on real data (Real-data Households) and those that focus on market information (SMFG households). We simulate the economic operations at various proportions of SMFG households: 0%, 25%, 50%, 75%, and 100%. Figure 5 displays the microeconomic indicators and macroeconomic GDP, with each subplot illustrating the average values of different economic indicators for different households (bars in left Y-axis) and their aggregate values (dots in right Y-axis). GDP, being a macroeconomic indicator, did not have corresponding values plotted for households.
The results indicate that, across all proportions, the microeconomic indicators of SMFG households, such as utility, wealth, and income, are consistently higher than Real-data households. As the proportion of SMFG households increases, there is an upward trend in economic indicators. This suggests that the presence of SMFG households positively impacts both individual and overall economic development, and our SMFG method also works when some households do not meet the assumption of focusing on market information. Therefore, governments should encourage households to pay more attention to market information in decision-making.
@@ 6 Conclusion
This paper innovatively models the optimal macroeconomic policy problem based on the Stackelberg mean field game, and proposes a solution method involving pre-training based on real data and a model-free Stackelberg mean field reinforcement learning algorithm. Our method addresses the limitations of existing approaches, including dynamic interactions and feedback between governments and households, interactions among large populations of households, and solving complex tasks without relying on environmental transitions. Experimental results have showcased the performance of the SMFG method. In conclusion, this paper contributes an effective approach to modeling and solving optimal macroeconomic policies in the field of AI for economics and AI for social good.
Impact Statements
This paper presents work whose goal is to advance the fields of AI for social good and AI for economics. Our work aims to offer suggestions and references for governments and the people, yet it must not be rashly applied to the real world. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Economic Model Details
$S(W^{*}_{t})=D(W^{*}_{t}),\mathcal{L}_{t}=\sum^{N}_{i}e_{t}^{i}h_{t}^{i}.$
$K_{t+1}+B_{t+1}-A_{t+1}=\left(r_{t}+1\right)(K_{t}+B_{t}-A_{t})$
$Y_{t}=C_{t}+G_{t}+X_{t}$
@@ Appendix A Economic Model Details
  Initialize $\tilde{Q}_{\phi^{l}},\tilde{Q}_{\phi^{l}_{-}},\tilde{Q}_{\phi^{f}},\tilde{Q}_{\phi^{f}_{-}},\pi_{\theta^{l}},\pi_{\theta^{l}_{-}},\pi_{\theta^{f}},\pi_{\theta^{f}_{-}}$.
@@
     Recieve initial observation $\mathbf{s}_{t}=\{s^{l}_{t},s^{f1}_{t},...,s^{fN}_{t}\}$.
@@
        For leader agent, select action $a^{l}_{t}=\pi_{\theta^{l}}(s^{l}_{t})+\mathcal{N}_{t}$; for each follower agent $i$, select action $a^{fi}=\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})+\mathcal{N}_{t}$ w.r.t. the current policy and exploration.
        Execute actions $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\}$ and observe reward $\mathbf{r}_{t}=\{r_{t}^{l},r^{f1}_{t},...,r^{fN}_{t}\}$ and next observation $\mathbf{s}_{t+1}$.
@@
           For the leader agent, set
$y^{l}_{t}=r_{t}^{l}+\gamma\tilde{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})}.$
@@
$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right]$
@@
$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}\tilde{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right].$
           For the representative follower agent, set
$y^{f}_{t}=r^{f}_{t}+\gamma\tilde{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}).$
           Update follower’s critic by minimizing the loss:
$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{f}}\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}.$
           Update follower’s actor by policy gradient:
$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}\tilde{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]}.$
@@
        Update target network parameters for each agent with learning rates $\tau_{\phi}$ and $\tau_{\theta}$:
@@
Appendix B Additional Results
B.1 Baselines
Baselines for solving SMFGs
Current research on model-based methods for SMFG is limited by strong assumptions and simplifications, rendering them ineffective for complex problem-solving. However, there is a lack of model-free algorithms for solving SMFG in continuous decision spaces. Therefore, this study opts to compare with popular reinforcement learning algorithms, including DDPG, MADDPG, IDDPG, and mean field MARL. Table 2 provides detailed information on the algorithms used by the leader and follower agent in these baselines.
Table 2: The details of the algorithms used by the leader and follower agents in the baselines for solving SMFG.
Baselines Follower’s Algorithm Leader’s Algorithm
Rule-based Random/Behavior cloning Rule-based/Free market
DDPG Random/Behavior cloning DDPG
MADDPG MADDPG MADDPG
IDDPG IDDPG IDDPG
MF-MARL Mean Field MARL DDPG
SMFRL SMFRL SMFRL
Baselines for economic policies
We compare 5 economic policies with our method SMFG:
(1) Free Market: A market without policy intervention.
(2) Saez Tax (Saez, 2001): The Saex tax policy is often considered a suggestion for specific tax reforms in the real world. Details are shown in Appendix C.
(3) U.S. Federal Tax: Real data from OECD  [^3] https://data.oecd.org/ in 2022 for this policy.
(4) AI Economist (Zheng et al., ): This is a two-level MARL method based on Proximal Policy Optimization (PPO). In the first phase, households’ policies are trained from scratch in a free-market (no-tax) environment. In the second phase, households continue to learn under an RL social planner.
(5) AI Economist-BC: For fairness in comparison, we evaluated the AI Economist method with behavior cloning as pre-training to determine its effectiveness.
B.2 Efficiency-Equity Tradeoff
In Figure 6, we also present the efficiency (per capita GDP) and equity (wealth Gini) corresponding to various algorithms for solving SMFGs. These methods result in similar Gini indices, with SMFRL achieving the highest per capita GDP. Apart from a strong emphasis on societal fairness, SMFRL emerges as the algorithm that can best strike a balance between efficiency and equity.
Figure 6: Comparison of macroeconomic indicators under different algorithms for solving SMFGs.
Table 3: Test results for 6 tax policies on 5 macroeconomic indicators ($N=100$ $N=1000$).
Algorithms $N=100$ $N=1000$

### v2

@@ 5 Experiment
In this section, we conduct experiments within the TaxAI environment to answer the following three questions: (1) How effective is our SMFRL algorithm in solving dynamic SMFGs? (2) For macroeconomic policy-making issues, how does our method compare to real-world strategies, economic models, and existing AI-based methods? (3) Given the likely real-world scenario where some households do not adopt the SMFG household policy learned by our method, does our approach remain effective?
TaxAI Environment
TaxAI [41] is one of the rare reinforcement learning environments designed to study optimal macroeconomic policies. It simulates dynamic interactions between large-scale households and the government, making it particularly well-suited for experimenting with SMFGs. TaxAI models the optimal tax policy problems where the government optimizes tax policies to promote economic growth (i.e. GDP) and social equity (i.e. Gini). In response, households optimize their labor supply and savings strategies to maximize their long-term utility.
5.1 Solving Dynamic SMFGs
Metrics and Baselines
We compare several classic multi-agent reinforcement learning (MARL) algorithms based on three metrics: leader’s payoff, exploitability, and social welfare, as defined in Section 3.3. Optimal Value: The ideal value of exploitability is 00, indicating that the follower’s policy achieves the best response to the leader’s policy. The upper bound of Social Welfare is $N$, given that household utility in TaxAI cannot exceed 1. For Leader’s Payoff, no maximum has been established. MF-MARL [61]: Utilizes the DDPG [36] algorithm for the leader and a mean-field MARL approach for the followers. MADDPG [38] and IDDPG [17]: Both feature shared actor and critic networks among followers. DDPG+Real: DDPG for the leader and behavior cloning [57] from real data for followers. U.S. policy+Real: U.S. federal tax policy for the leader with behavior cloning for the followers.
Results Analysis
Table 1 presents the performance metrics for SMFRL and five baselines, alongside their optimal values, averaged over 10 evaluations of models trained with at least five seeds. SMFRL notably outperforms the baselines in achieving higher leader’s payoff, demonstrating its effectiveness in optimizing the leader’s objective. The DDPG+Real method, using a real-data trained policy, attains a substantial payoff but exhibits a significant gap from the households’ optimal response, as indicated by its higher exploitability. In contrast, SMFRL significantly reduces exploitability for both $N=100$ and $N=1000$, suggesting that its policy for households closely approximates the best response to the leader’s strategy, which is crucial for solving Stackelberg mean field games. Moreover, SMFRL closely approaches the optimal social welfare value, indicating that its solutions are the most socially optimal among the considered baselines. In summary, the SMFRL algorithm effectively learns policies that maximize the leader’s objective while ensuring that followers attain the best response to the leader’s policy, closely approaching the social optimum. The training curves are presented in Figure 4, and the test results for macroeconomic indicators are detailed in Table 1 in Appendix D.
5.2 Macroeconomic Policy-Making
Table 2: Test results 6 tax policies across five macroeconomic indicators at two numbers of followers, $N=100$ and $N=1000$.
Households’ Policies Macroeconomic Indicators
@@ 5.2 Macroeconomic Policy-Making
Number Per Capita GDP Social Welfare Income Gini Wealth Gini Years
$\mathbf{N=100}$ Free Market 1.37e+05 32.97 0.89 0.92 1.10
Saez Tax 2.34e+12 73.82 0.21 0.38 300.00
US Federal Tax 4.88e+11 94.19 0.40 0.40 289.55
AI Economist-BC 4.24e+12 97.24 0.54 0.52 299.55
AI Economist 1.26e+05 72.81 0.88 0.91 1.00
SMFG 9.59e+12 96.87 0.52 0.51 300.0
$\mathbf{N=1000}$ Free Market 1.41e+05 334.79 0.90 0.93 1.00
Saez Tax 6.35e+11 498.88 0.68 0.73 100.578
US Federal Tax 1.41e+05 351.17 0.89 0.93 1.00
AI Economist-BC N/A N/A N/A N/A N/A
AI Economist N/A N/A N/A N/A N/A
SMFG 1.10e+13 968.94 0.54 0.53 300.00
Metrics and Baselines
In addressing macroeconomic policymaking, we utilize classical macroeconomic indicators from the real world to evaluate the performance of different policies. Per Capita GDP indicates the level of economic development, while Income Gini and Wealth Gini measure the inequality in households’ income and wealth, respectively; a lower Gini index suggests greater social equality. The Years metric represents the sustainable duration of an economy, capped at 300 years; economies may terminate earlier due to excessive inequality or inadequate productivity to meet the populace’s needs. We compare several methods for solving macroeconomic policies: the Free Market [5] policy entails no government intervention; the Saez Tax [50] represents a well-known economic approach, often advocated for specific tax reforms in practice. The U.S. Federal Tax is the real policy from the OECD data  [^2] https://data.oecd.org/. In these approaches, households employ strategies based on behavior cloning from real data. Additionally, the AI Economist [62] represents an AI-based policy using a two-level Proximal Policy Optimization (PPO) framework. The AI Economist-BC method leverages behavior cloning policy on real-world data for pretraining.
Results Analysis
Table 2 presents the average values of macroeconomic indicators across multiple test rounds for different algorithms under varying numbers of households. The results reveal that the Saez Tax method achieves the lowest income Gini and wealth Gini when the number of households is small ($N=100$). However, our SMFG method yields higher per capita GDP and social welfare. At $N=1000$, the advantages of the SMFG algorithm become even more pronounced compared to other methods. It secures the longest duration (Years), the highest GDP, and the lowest income Gini and wealth Gini. The social welfare derived from the SMFG approach closely approaches the optimal value of 1000. This demonstrates that our method can learn the macroeconomic policy that simultaneously optimizes economic growth, social equity, and social welfare, surpassing existing economic approaches like the Saez Tax and AI-based policies (i.e. AI Economist). We provide the corresponding training curves in Figure 5 in Appendix D.
5.3 Effectiveness Analysis in Diverse Households
Figure 3: The left subfigure illustrates a scenario with heterogeneous households adopting various policies. The right subfigure presents test results for households’ decisions (consumption, labor), microeconomic indicators (utility, wealth, income), and macroeconomic indicators (GDP) when there are different proportions of SMFG households ($N=100$).
Problem Description
The above experiments have already demonstrated the effectiveness of our method in addressing macroeconomic policy-making problems. However, when extending SMFG approach to real-world scenarios, where governments and households implement leader and follower policies trained via the SMFG method (referred to as SMFG government and SMFG households), there remain partial households that might not adopt the SMFG follower policy. In economic terms, this scenario can be interpreted as the government introducing new policies and providing recommendations to households, some of whom may choose not to follow these recommendations. When faced with heterogeneous households that adopt varying strategies, the question arises: is our SMFG method still effective? In experiments, to model households that do not adopt the SMFG policy, referred to as Real-data Households, we utilize behavior cloning policy trained by real data. Related experimental details are shown in Appendix D.4. We set different proportions of SMFG households (0%, 25%, 50%, 75%, and 100%) and evaluate our SMFG method using microeconomic and macroeconomic indicators. In Figure 3, each subplot displays average values for two types of households (bars on the left Y-axis) and their aggregate values (dots on the right Y-axis). GDP, a macroeconomic indicator, is shown only in aggregate, not for individual households.
Result Analysis
From Figure 3, we identify several key findings: (1) SMFG Households outperform Real-data Households across microeconomic indicators such as wealth, income, and utility, which should encourage more households to adopt the SMFG policy; (2) The presence of heterogeneous households does not prevent SMFG Households from achieving near-optimal utility, capped at 1 in the TaxAI environment; (3) The lower labor contribution from Real-data Households leads to a decrease in Per Capita GDP as their proportion increases. Thus, the more households adopt the SMFG follower policy, the more effective the SMFG method becomes.
@@ 6 Conclusion
This paper innovatively models macroeconomic policymaking as a dynamic Stackelberg mean field game, with the government acting as the leader and large-scale households as followers. We introduce a data-driven Stackelberg mean field reinforcement learning algorithm that effectively solves SMFGs. When applied to macroeconomic policy formulation, our method significantly outperforms existing economic approaches and AI-based policies, especially in scenarios involving large numbers of households. Furthermore, we demonstrate the robustness of our approach, showing that it retains effectiveness even when some households do not adopt the SMFG policy. In conclusion, this paper contributes an effective approach to modeling and solving optimal macroeconomic policies in the field of AI for economics and AI for social good.
Appendix A SMFRL algorithm pseudocode
@@
  Initialize ${Q}_{\phi^{l}},{Q}_{\phi^{l}_{-}},{Q}_{\phi^{f}},{Q}_{\phi^{f}_{-}},\pi_{\theta^{l}},\pi_{\theta^{l}_{-}},\pi_{\theta^{f}},\pi_{\theta^{f}_{-}}$.
@@
     Receive initial state $\mathbf{s}_{t}=\{s^{l}_{t},s^{f1}_{t},...,s^{fN}_{t}\}$.
@@
        For the leader agent, select action $a^{l}_{t}=\pi_{\theta^{l}}(s^{l}_{t})+\mathcal{N}_{t}$; for each follower agent $i$, select action $a^{fi}=\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})+\mathcal{N}_{t}$ w.r.t. the current policy and exploration.
        Execute actions $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\}$ and observe reward $\mathbf{r}_{t}=\{r_{t}^{l},r^{f1}_{t},...,r^{fN}_{t}\}$ and next state $\mathbf{s}_{t+1}$.
@@
           For each follower agent, perform inner updates:
           for k = 1 to inner-update-cycles do
              Set
$y^{f}_{t}=r^{f}_{t}+\gamma{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}).$
              Update follower’s critic by minimizing the loss:
$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{f}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}.$
              Update follower’s actor by policy gradient:
$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]}.$
           end for
           After follower updates, update the leader’s networks:
           Set
$y^{l}_{t}=r_{t}^{l}+\gamma{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})}.$
@@
$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right]$
@@
$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right].$
@@
        Update target network parameters:
@@
Appendix B Assumptions and Limitations
Assumptions
This paper models the problem of macroeconomic policy-making as a Dynamic Stackelberg Mean Field Game, based on the following assumptions: (1) Homogeneous followers: We assume that a large-scale group of households is homogeneous. They can use different characteristics as observations to influence decisions, but there are commonalities in human behavioral strategies. (2) Rational Expectations: We assume that both macro and micro agents engage in rational decision-making, adjusting their future expectations based on observed information. However, in reality, the level of rationality varies among different households. Most households exhibit bounded rationality, and their expectations and preferences differ accordingly. (3) Experimental environment: We validate our approach through experiments in the TaxAI environment, based on the assumption that results within this environment can provide insights applicable to real-world scenarios. Addressing and potentially relaxing these assumptions will be a primary focus of our future research.
Limitations
The limitations of our SMFG method will be thoroughly investigated in future work: (1) We plan to consider dynamic games involving multiple leaders and large-scale followers to explore policy coordination across various macroeconomic sectors. (2) We will continue to develop theoretical proofs for the equilibrium solutions in Stackelberg mean field games. Currently, our approach is empirically demonstrated by showing that followers converge toward their best responses and that the leader achieves higher performance compared to other baselines. (3) We intend to examine dynamic games between a leader and a large, heterogeneous group of followers, including scenarios where followers dynamically alter their strategies, to determine the optimal leader policy. Addressing these limitations will provide further insights applicable to real-world scenarios.
Appendix C Impact Statements
This paper presents work whose goal is to advance the fields of AI for social science and AI for economics. Our work aims to offer suggestions and references for governments and the people, yet it must not be rashly applied to the real world. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix D Additional Results
Compute Resources
All experiments are run on 2 workstations: A 64-bit server with dual AMD EPYC 7742 64-Core Processors @2.25 GHz, 256 cores, 512 threads, 503GB RAM, and 2 NVIDIA A100-PCIE-40GB GPU. A 64-bit workstation with Intel Core i9-10920X CPU @ 3.50GHz, 24 cores, 48 threads, 125 GB RAM, and 2 NVIDIA RTX2080 Ti GPUs. The following Table 3 shows the approximate training times for several algorithms.
Table 3: the approximate training times for several algorithms for solving dynamic SMFGs.
Algorithm N=100 (hours) N=1000 (hours)
SMFRL 4 14
MADDPG 3 16
MF-MARL 4 9
IDDPG 2 6
DDPG+Real 1 5
Rule-based+Real 0.25 2
D.1 Training Curves for Dynamic SMFGs
Due to the challenging nature of the task, the baselines we compared did not perform well in solving dynamic SMFGs. Therefore, we utilized behavior cloning as a pre-training strategy for the follower agents. Therefore, we compared SMFRL with 5 baselines in 4 different experimental setups: without behavior cloning as pre-training for follower agents at N=100 and N=1000 (marked as N=100 without BC and N=1000 without BC); with BC-based pre-training for follower agents at N=100 and N=1000 (N=100-BC; N=1000-BC). Figure 4 illustrates the training curves of 4 key macroeconomic indicators under these three settings. The solid line represents the average value of the metrics across the 5 random seeds, while the shaded area represents the standard deviation. Each row corresponds to one setting, and each column to a macroeconomic indicator, including per capita GDP, social welfare, income Gini, and wealth Gini. A rise in per capita GDP indicates economic growth, an increase in social welfare implies happier households and a lower Gini index indicates a fairer society. Each subplot’s Y-axis represents the indicators’ values, and the X-axis represents the training steps. Table 4 displays the test results of the 6 algorithms across 5 indicators, with each column corresponding to an experimental setting.
Table 4: Performance metrics of six algorithms across different settings on key economic indicators.